methane(CH⁴) burns in oxygen to produce carbon dioxide and water vapor , what is the number of(CH²) MOLECULES produced when 3.2 L of oxygen are comsumed?

1
Expert's answer
2021-11-22T17:19:08-0500

Methane(CH4) burns in oxygen to produce carbon dioxide and water vapor.

What is the number of (CO2) molecules produced when 3.2 L of oxygen are comsumed?


Solution:

Calculate the moles of oxygen gas (O2):

At STP, one mole of any gas occupies a volume of 22.4 L.

Hence,

(3.2 L O2) × (1 mol O2 / 22.4 L O2) = 0.143 mol O2


Balanced chemical equation:

CH4(g) + 2O2(g) → CO2(g) + 2H2O(g)

According to stoichiometry:

2 mol of O2 produces 1 mol of CO2

Thus, 0.143 mol of O2 produces:

(0.143 mol O2) × (1 mol CO2 / 2 mol O2) = 0.0715 mol CO2


One mole of any substance contains 6.022×1023 particles.

Hence,

Number of CO2 molecules = (0.0715 mol) × (6.022×1023 molecules / 1 mol) = 4.3×1022 molecules

Number of CO2 molecules = 4.3×1022 molecules


Answer: 4.3×1022 molecules of CO2 molecules produced